# Environment as Policy: Learning to Race in Unseen Tracks

### v2

@@ I Introduction
Reinforcement learning (RL) involves agents learning through trial and error by interacting with a pre-defined environment and maximizing the rewards based on these interactions. It has proven highly effective in various robotic control applications, demonstrating remarkable task performance across scenarios like dexterous manipulation [1, 2, 3, 4], quadrupedal locomotion [5, 6], and agile quadrotor flight [7, 8, 9, 10, 11]. Previous research has shown that RL can even outperform human champions in the task of drone racing [12, 13], where an RL agent flies a quadcopter drone through a complex track at high speed, requiring quick decision-making and precise control.
However, while RL agents excel within the specific distributions they are trained on, they struggle with out-of-distribution configurations and may require retraining from scratch for even minor configuration changes [14]. To improve adaptability and generalization, extending the RL framework to train and perform across a broader range of distributions is essential, enabling agents to handle more diverse and dynamic environments effectively. However, training directly on a wide distribution of scenarios can significantly degrade the efficiency and effectiveness of the learning process [15]. As the distribution broadens, the complexity of policy exploration increases, making it harder for the agent to identify and implement effective solutions.
Domain randomization is a commonly used technique to improve the learning capability of RL agents across a broader range of tasks [14, 16, 17, 18]. This approach varies the parameters of the training environment to expose the agent to a wide variety of potential deployment scenarios. By learning in diverse conditions, the agent develops robust policies less likely to overfit the specific characteristics of a single environment. Domain randomization is often combined with curriculum learning, where the complexity and variability of training scenarios are incrementally increased [18]. By starting with less challenging environments and gradually introducing greater variability, the agent learns to perform well in a wide set of scenarios without becoming overwhelmed during the early stages of training. However, curriculum learning often depends on manual design, introducing human biases. This reliance on fixed, non-adaptive progressions can limit training diversity and restrict the agent’s ability to generalize to new, real-world situations [19]. These effects are especially prominent in the challenging task of drone racing, where the platform’s high agility and the need for rapid decision-making amplify the difficulties. Therefore, the challenge of agile drone racing on an unseen race track in real-world conditions remains largely unexplored.
@@ II Related Works
Reinforcement Learning for Robotics Reinforcement Learning (RL) has been extensively applied in robotics to facilitate continuous control in complex and dynamic environments. Recent studies have achieved remarkable success in various robotic domains. These include quadrotor control [12, 13, 20, 9, 21, 22], legged locomotion [23, 6], and manipulation [24, 25], showcasing significant advancements in the field.
Autonomous Drone Racing Autonomous drone racing is an emerging field in which drones must navigate through predefined waypoints in the shortest possible time. Success in this field requires integrating advanced hardware and sophisticated algorithms capable of perceiving the environment, planning optimal paths, and executing actions in real-time [26]. [13] demonstrated that RL controllers can outperform traditional optimization-based approaches, such as those proposed in [27, 28], for autonomous drone racing. This advantage arises because RL can effectively handle highly non-linear dynamical systems and complex objectives, which pose significant challenges for conventional optimization methods like Model Predictive Control (MPC) [29, 30]. Recent studies have shown that policies trained with deep reinforcement learning (DRL) can surpass even human world champions, employing vision-based state estimation for the RL controller. Furthermore,  [31, 21] have demonstrated using RL to learn drone racing from image inputs directly.
Environment Shaping for Reinforcement Learning Environment shaping has been widely applied in reinforcement learning by designing a distribution of environments that progressively improves the agent’s performance. Previous works like [32, 33, 34, 35] on environment shaping have been successful in simulation tasks like Bipedal Walker and maze games, where environments are easily parameterized for manual curriculum design. However, this process becomes more challenging for more complex real-world robotic tasks. Domain randomization is a common solution, where environments are sampled from a predefined range to enable robust task execution. For instance,  [18, 17] used domain randomization to achieve dexterous manipulation in the real world. Additionally,  [5] proposed an adaptive terrain curriculum using a particle filter to sample environment parameters, enabling quadrupedal locomotion across various terrains in real-world scenarios.
@@ III-A Racing Policy Training
The autonomous racing task can be framed as an optimization problem, where the objective is to minimize the time it takes for an agile quadrotor to pass through a predefined sequence of gates [26], as illustrated in Fig. 2. In this task, we define the observations as $\bm{o}_{\mathrm{racing}}=\left[\tilde{\bm{R}},\bm{v},\bm{\omega},a_{\text{prev}},\delta\bm{p}_{1},\delta\bm{p}_{2}\right]$, where $\tilde{\bm{R}}\in\mathbb{R}^{6}$ is a vector comprising the first two columns of $\bm{R}_{WB}$ [36], $\bm{v}\in\mathbb{R}^{3}$ and $\bm{\omega}\in\mathbb{R}^{3}$ denote the linear and angular velocity of the drone, $a_{\text{prev}}$ represents the previous action from the actor policy, and $\delta\bm{p}_{1},\delta\bm{p}_{2}\in\mathbb{R}^{12}$ represent the relative difference in position of the four next gate corners ($4\times 3$) in the world frame. Here, $\delta\bm{p}_{1}$ represents the difference of the four corners of the next gate to pass between the current quadrotor position, and $\delta\bm{p}_{2}$ represents the positional difference of the corners between the next gate to pass and the gate after the next gate to pass on the race track. The total reward at time $t$, denoted as $r_{t}$, consists of several components:
@@
where $r_{t}^{\mathrm{prog}}$ represents progress toward passing the next gate [37], $r_{t}^{\mathrm{act}}$ penalizes changes in actions from the previous time step, $r_{t}^{\mathrm{br}}$ discourages high body rates to ensure a stable flying behavior, $r_{t}^{\mathrm{pass}}$ is a binary reward for successfully passing the next gate, and $r_{t}^{\mathrm{crash}}$ is a binary penalty applied when a collision occurs, which also terminates the episode. The reward components are formulated as follows
@@
[…]
@@ III-B1 MDP Formulation
$\vspace{-0.8em}r^{env}_{t}=\begin{cases}R\frac{N_{\mathrm{env}}-\mathrm{rank}}{N_{\mathrm{env}-r_{\mathrm{upper}}}}&\text{if }rank>r_{\mathrm{upper}}\\ R&\text{if }rank\in[r_{\mathrm{lower}},r_{\mathrm{upper}}]\\ R\frac{\mathrm{rank}}{r_{\mathrm{lower}}}&\text{if }rank<r_{\mathrm{lower}}\end{cases}\vspace{0.3em}$ (3)
@@ III-B2 Environment Shaping
To accelerate data collection, we train our environment agent and the racing agent using multiple parallel simulation environments. To handle varying numbers of environments, the environment policy independently updates the track layout for each environment. Since all environments initially share the same track layout, the racing policy experiences a gradual increase in layout complexity, which facilitates smoother learning during the early stages of training. We use the notation $\mathrm{Env}_{i}\leftarrow\mathrm{Env}_{i-1}\oplus a_{i-1}$ to represent changes in the individual environment. For each new environment, the environment policy is trained for every $N_{\text{freq}}$ racing policy update, where $N_{\text{freq}}$ determines how much more frequent should $\pi_{\mathrm{racing}}$ than $\pi_{\mathrm{env}}$. Since $N_{\text{freq}}$ is typically much greater than 1 (i.e., $N_{\text{freq}}\gg 1$), the environment policy is updated less frequently than the racing policy. To handle this, we choose the off-policy RL algorithm Soft Actor-Critic (SAC) [38] to update the parameters of the environment policy. The complete framework containing the environment policy and the racing policy during the training is presented in Algorithm 1.
@@ IV-A Experimental Setup
TABLE I: We compare the success rate (SR) and lap time (LT) of our method against four baselines. Six different unseen racetracks are evaluated in a realistic BEM simulation [39], including three 2D tracks and three 3D tracks. Here apart from the our and baseline approaches, we also include the Single-track RL as a reference, which is a vanilla PPO agent trained and tested on individual track.
Our training framework is built on the Flightmare simulator [40], with reward functions and PPO hyperparameters aligned with methods from previous research [12]. For our environment policy, we implemented a vectorized Gym [41] environment with the same number of environments as the agent, using a total of 100 environments. The environment policy uses the SAC implementation from Stable Baselines 3 [42], with fine-tuned parameters for our task setting. Specifically, we set the gradient steps of SAC to 100 to accelerate the training of the environment policy and the entropy coefficient to 1.0 for better exploration.
During the training process, the racing policy is updated at every iteration, while the environment policy is updated every 100 iterations. We sometimes encounter environments getting stuck on infeasible tracks during training. To address this, if the evaluation SR stays at zero for three consecutive runs, we reset the environment to the initial race track. We selected an 8-gate oval track for the initial layout due to its simplicity, making it easier for the racing policy to learn and adapt quickly. For the environment policy’s action space, we constrained the gate movements as follows: along the x and y axes, the range is $[-1.0,1.0]$$\mathrm{m}$; along the z-axis, the range is $[-0.2,0.2]$$\mathrm{m}$, and for the yaw, the range is $[-\frac{\pi}{30},\frac{\pi}{30}]$ $\mathrm{rad}$. For the relative ranking range, we set $r_{lower}$ to the 50th percentile of all environments and $r_{upper}$ to the 90th percentile. These hyperparameters are empirically found to lead to the best performance during the ablation experiments. Our method is trained for 800M data samples, which is equivalent to 12 hours of training using an NVIDIA TITAN Xp GPU.
To evaluate the performance of our approach, we tested the racing policy on several fixed racetracks with varying difficulties and complexities. We selected three 2D tracks, where all gates are positioned at the same height along the z-axis, including Figure 8, Kidney, and 2D Big S. Additionally, we chose three 3D tracks, where the gates have varying heights along the z-axis, including 3D Figure 8, 3D Big S, and Twist, as shown in Fig. 2. All of these tracks are significantly different than the initial training track, with most having a different number of gates than during training. For the evaluation, we primarily used two metrics: success rate (SR %) and lap time (LT [$\mathrm{s}$]). The success rate is calculated as the ratio of successful runs, where the drone passes all the gates without crashing, over the total number of trials. Lap time refers to the drone’s total time to successfully complete a race track. These metrics are commonly used in drone racing and are essential in evaluating whether the policy enables fast and stable flight performance [31, 21].
@@
To evaluate the generalization ability of our proposed methods, we compared them with three baseline methods: (i) RL policy without curriculum: This method shares the same initial environment setup as ours but does not use a curriculum for training. (ii) Domain randomization from [7]: In this approach, the initial environment is the same as in our method. However, environment policy actions are sampled randomly within the action space; we use the identical configurations from [7]. (iii) Particle Filter Curriculum from [5]: This method also uses the same initial environment as ours but relies on a particle filter to sample environments.
@@
Range [$\mathrm{m}$] Methods SR [%] LT [s]

### v3

@@ I Introduction
Reinforcement learning (RL) involves agents learning through trial and error by interacting with a pre-defined environment and maximizing the rewards based on these interactions. It has proven highly effective in various robotic control applications, demonstrating remarkable task performance across scenarios like dexterous manipulation [andrychowicz2020learning, kalashnikov2018scalable, kalashnikov2022scaling, aljalbout2024role], quadrupedal locomotion [lee2020learning, kumar2021rma], and agile quadrotor flight [song2021autonomous, messikommer2024contrastive, xing2024contrastive, krinner2025accelerating, romero2024actor2]. Previous research has shown that RL can even outperform human champions in the task of drone racing [kaufmann23champion, Song23Reaching], where an RL agent flies a quadcopter drone through a complex track at high speed, requiring quick decision-making and precise control.
However, while RL agents excel within the specific distributions they are trained on, they struggle with out-of-distribution configurations and may require retraining from scratch for even minor configuration changes [DR]. To improve adaptability and generalization, extending the RL framework to train and perform across a broader range of distributions is essential, enabling agents to handle more diverse and dynamic environments effectively. However, training directly on a wide distribution of scenarios can significantly degrade the efficiency and effectiveness of the learning process [bengio2009curriculum]. As the distribution broadens, the complexity of policy exploration increases, making it harder for the agent to identify and implement effective solutions.
Domain randomization is a commonly used technique to improve the learning capability of RL agents across a broader range of tasks [DR, DR2, openai2019learning, openai2019solving]. This approach varies the parameters of the training environment to expose the agent to a wide variety of potential deployment scenarios. By learning in diverse conditions, the agent develops robust policies less likely to overfit the specific characteristics of a single environment. Domain randomization is often combined with curriculum learning, where the complexity and variability of training scenarios are incrementally increased [openai2019solving]. By starting with less challenging environments and gradually introducing greater variability, the agent learns to perform well in a wide set of scenarios without becoming overwhelmed during the early stages of training. However, curriculum learning often depends on manual design, introducing human biases. This reliance on fixed, non-adaptive progressions can limit training diversity and restrict the agent’s ability to generalize to new, real-world situations [kong2021adaptive]. These effects are especially prominent in the challenging task of drone racing, where the platform’s high agility and the need for rapid decision-making amplify the difficulties. Therefore, the challenge of agile drone racing on an unseen race track in real-world conditions remains largely unexplored.
@@ II Related Works
Reinforcement Learning for Robotics Reinforcement Learning (RL) has been extensively applied in robotics to facilitate continuous control in complex and dynamic environments. Recent studies have achieved remarkable success in various robotic domains. These include quadrotor control [kaufmann23champion, Song23Reaching, hwangbo2017control, xing2024contrastive, xing2024bootstrapping, messikommer2024student], legged locomotion [hwangbo2019learning, kumar2021rma], and manipulation [fu2023deep, gu2017deep], showcasing significant advancements in the field.
Autonomous Drone Racing Autonomous drone racing is an emerging field in which drones must navigate through predefined waypoints in the shortest possible time. Success in this field requires integrating advanced hardware and sophisticated algorithms capable of perceiving the environment, planning optimal paths, and executing actions in real-time [hanover2024autonomous]. [Song23Reaching] demonstrated that RL controllers can outperform traditional optimization-based approaches, such as those proposed in [romero2022model, romero2024actor], for autonomous drone racing. This advantage arises because RL can effectively handle highly non-linear dynamical systems and complex objectives, which pose significant challenges for conventional optimization methods like Model Predictive Control (MPC) [falanga2018pampc, xing2023autonomous]. Recent studies have shown that policies trained with deep reinforcement learning (DRL) can surpass even human world champions, employing vision-based state estimation for the RL controller. Furthermore,  [geles2024demonstrating, xing2024bootstrapping] have demonstrated using RL to learn drone racing from image inputs directly.
Environment Shaping for Reinforcement Learning Environment shaping has been widely applied in reinforcement learning by designing a distribution of environments that progressively improves the agent’s performance. Previous works like [portelas2019, parkerholder2023, jiang2022, dennis2021] on environment shaping have been successful in simulation tasks like Bipedal Walker and maze games, where environments are easily parameterized for manual curriculum design. However, this process becomes more challenging for more complex real-world robotic tasks. Domain randomization is a common solution, where environments are sampled from a predefined range to enable robust task execution. For instance,  [openai2019solving, openai2019learning] used domain randomization to achieve dexterous manipulation in the real world. Additionally,  [lee2020learning] proposed an adaptive terrain curriculum using a particle filter to sample environment parameters, enabling quadrupedal locomotion across various terrains in real-world scenarios.
@@ III-A Racing Policy Training
The autonomous racing task can be framed as an optimization problem, where the objective is to minimize the time it takes for an agile quadrotor to pass through a predefined sequence of gates [hanover2024autonomous], as illustrated in Fig. 2. In this task, we define the observations as $\bm{o}_{\mathrm{racing}}=\left[\tilde{\bm{R}},\bm{v},\bm{\omega},a_{\text{prev}},\delta\bm{p}_{1},\delta\bm{p}_{2}\right]$, where $\tilde{\bm{R}}\in\mathbb{R}^{6}$ is a vector comprising the first two columns of $\bm{R}_{WB}$ [zhou2019continuity], $\bm{v}\in\mathbb{R}^{3}$ and $\bm{\omega}\in\mathbb{R}^{3}$ denote the linear and angular velocity of the drone, $a_{\text{prev}}$ represents the previous action from the actor policy, and $\delta\bm{p}_{1},\delta\bm{p}_{2}\in\mathbb{R}^{12}$ represent the relative difference in position of the four next gate corners ($4\times 3$) in the world frame. Here, $\delta\bm{p}_{1}$ represents the difference of the four corners of the next gate to pass between the current quadrotor position, and $\delta\bm{p}_{2}$ represents the positional difference of the corners between the next gate to pass and the gate after the next gate to pass on the race track. The total reward at time $t$, denoted as $r_{t}$, consists of several components:
@@
where $r_{t}^{\mathrm{prog}}$ represents progress toward passing the next gate [yunlong2021racing], $r_{t}^{\mathrm{act}}$ penalizes changes in actions from the previous time step, $r_{t}^{\mathrm{br}}$ discourages high body rates to ensure a stable flying behavior, $r_{t}^{\mathrm{pass}}$ is a binary reward for successfully passing the next gate, and $r_{t}^{\mathrm{crash}}$ is a binary penalty applied when a collision occurs, which also terminates the episode. The reward components are formulated as follows
@@
$r_{t}^{\mathrm{crash}}$ $=\alpha_{5}\quad\text{if robot crashes (gates, ground)}.$
@@ III-B1 MDP Formulation
$\vskip-8.00003ptr^{env}_{t}=\begin{cases}R\frac{N_{\mathrm{env}}-\mathrm{rank}}{N_{\mathrm{env}-r_{\mathrm{upper}}}}&\text{if }rank>r_{\mathrm{upper}}\\ R&\text{if }rank\in[r_{\mathrm{lower}},r_{\mathrm{upper}}]\\ R\frac{\mathrm{rank}}{r_{\mathrm{lower}}}&\text{if }rank<r_{\mathrm{lower}}\end{cases}\vskip 3.00003pt$ (3)
@@ III-B2 Environment Shaping
To accelerate data collection, we train our environment agent and the racing agent using multiple parallel simulation environments. To handle varying numbers of environments, the environment policy independently updates the track layout for each environment. Since all environments initially share the same track layout, the racing policy experiences a gradual increase in layout complexity, which facilitates smoother learning during the early stages of training. We use the notation $\mathrm{Env}_{i}\leftarrow\mathrm{Env}_{i-1}\oplus a_{i-1}$ to represent changes in the individual environment. For each new environment, the environment policy is trained for every $N_{\text{freq}}$ racing policy update, where $N_{\text{freq}}$ determines how much more frequent should $\pi_{\mathrm{racing}}$ than $\pi_{\mathrm{env}}$. Since $N_{\text{freq}}$ is typically much greater than 1 (i.e., $N_{\text{freq}}\gg 1$), the environment policy is updated less frequently than the racing policy. To handle this, we choose the off-policy RL algorithm Soft Actor-Critic (SAC) [SAC] to update the parameters of the environment policy. The complete framework containing the environment policy and the racing policy during the training is presented in Algorithm 1.
@@ IV-A Experimental Setup
TABLE I: We compare the success rate (SR) and lap time (LT) of our method against four baselines. Six different unseen racetracks are evaluated in a realistic BEM simulation [bauersfeld2021neurobem], including three 2D tracks and three 3D tracks. Here apart from the our and baseline approaches, we also include the Single-track RL as a reference, which is a vanilla PPO agent trained and tested on individual track.
Our training framework is built on the Flightmare simulator [song2020flightmare], with reward functions and PPO hyperparameters aligned with methods from previous research [kaufmann23champion]. For our environment policy, we implemented a vectorized Gym [GYM] environment with the same number of environments as the agent, using a total of 100 environments. The environment policy uses the SAC implementation from Stable Baselines 3 [stable-baselines3], with fine-tuned parameters for our task setting. Specifically, we set the gradient steps of SAC to 100 to accelerate the training of the environment policy and the entropy coefficient to 1.0 for better exploration.
[…]
To evaluate the performance of our approach, we tested the racing policy on several fixed racetracks with varying difficulties and complexities. We selected three 2D tracks, where all gates are positioned at the same height along the z-axis, including Figure 8, Kidney, and 2D Big S. Additionally, we chose three 3D tracks, where the gates have varying heights along the z-axis, including 3D Figure 8, 3D Big S, and Twist, as shown in Fig. 2. All of these tracks are significantly different than the initial training track, with most having a different number of gates than during training. For the evaluation, we primarily used two metrics: success rate (SR %) and lap time (LT $\mathrm{s}$]). The success rate is calculated as the ratio of successful runs, where the drone passes all the gates without crashing, over the total number of trials. Lap time refers to the drone’s total time to successfully complete a race track. These metrics are commonly used in drone racing and are essential in evaluating whether the policy enables fast and stable flight performance [geles2024demonstrating, xing2024bootstrapping].
@@
To evaluate the generalization ability of our proposed methods, we compared them with three baseline methods: (i) RL policy without curriculum: This method shares the same initial environment setup as ours but does not use a curriculum for training. (ii) Domain randomization from [song2021autonomous]: In this approach, the initial environment is the same as in our method. However, environment policy actions are sampled randomly within the action space; we use the identical configurations from [song2021autonomous]. (iii) Particle Filter Curriculum from [lee2020learning]: This method also uses the same initial environment as ours but relies on a particle filter to sample environments.
@@
Range $\mathrm{m}$] Methods SR [%] LT [s]